How many ways can a person toss a coin 8 times so that the number of heads is between 5 and 7 inclusive?
Answer #1

Combination formula :

combination of 8 choose 5 +8 choose 6 + 8 choose 7

=8!/(5!(8-5)!)+8!/(6!(8-6)!) + 8!/(7!(8-7)!)

=56+28+8

=92

A total of 2^8 = 256 possible flips

The probability is 92/256 = 0.3594